12. Compare and contrast the solubility of a solid and a gas in a liquid solvent. What factors influence solubility? What similarities and differences are there between gas and solid solubilities? How do the solubilities function in KMT terms? 13. a) Calculate the frequency and energy of 1 mol of photons with wavelength equal to 5.34x10m which are emitted by a compound. b) To what temperature would this quantity of energy raise 627 mL of water initially at 11.5 C? c) If one mole of atoms of a particular element requires 187 kJ of energy to liberate one electron per atom in the photoelectric effect, will this light provide sufficient energy? Why? 14. a) Write the equation for the reaction which occurs by combining manganese (III) chlorate with ammonium sulfide, if any occurs. b) Classify the reaction by type and what atoms do. c) What evidence of a reaction would one see if these two solutions were combined?
17. The reaction hydrochloric acid with sodium hydrogen carbonate produces carbon dioxide gas as one product. a) Write the equation for the reaction of these chemicals as aqueous solutions. 500. mL of equimolar concentration of each solution are combined in this reaction. b) If 2.461 daL of carbon dioxide gas are produced at 32.81 °C and 762.7 torr, constituting an 89.62% yield, what molarity of sodium hydrogen carbonate solution was originally reacted, presuming additive volumes? 18. If the molecules in a sample of gas weighing 58.62 g have an rms value of 3.37x10 m/s, at STP, a) what is the molar mass of the compound? b) What is its density at STP? c) If the aforementioned sample exists at 105.3°C and 758.6 torr, what is its volume?
Answer #1

#13:(a): Energy of single photon, E = h*c / \large \lambda = 6.626*10-34 J.s * 3*108 m/s / 5.34*10-7 m = 3.722*10-19 J / photon

1 mol photon contains 6.023*1023 photos

=> Energy of 1 mol photon = 6.023*1023 photos * (3.722*10-19 J / photon) = 2.242*105 J

(b): mass of 627 mL water = 627 mL * (1g / 1mL) = 627 g  

Initial temperature, Ti = 11.5 oC

Let the final temperature be "T oC"

Q = m*s*deltaT

=> 2.242*105 J = 627 g * 4.184 J/g.oC * (T - 11.5 oC)

=> (T - 11.5 oC) = 2.242*105 J / 627 g * 4.184 J/g.oC

=>  (T - 11.5 oC) = 85.5 oC

=> T = 85.5 oC + 11.5 oC = 97.0 oC (Answer)

(c): In (a), the energy liberated = 2.242*105 J * (1kJ/1000 J) = 224.2 kJ

Required energy = 187 kJ.

Since the liberated energy by the light 224.2 kJ is higher than 187 kJ, this light provides sufficient energy.

#14: (a): Manganese(III) chloride is Mn(ClO3)3

Ammonium sulfide is (NH4)2S

When Mn(ClO3)3 and (NH4)2S react, a precipitate of Mn2S3(s) formed.

Balanced equation: 2Mn(ClO3)3(aq) + 3(NH4)2S (aq) ----> Mn2S3(s) + 6NH4ClO3(aq)

(b):Since both the ions are displaces, the reaction type is double displacement reaction.

(c): Formation of a solid precipitate